OpenStudy (anonymous):

Vernon has obtained a $105,000, 5/1 30-year ARM at 5%. During the first 3 years, he has an option of paying interest only. If he accepts this offer, what would be his initial payment? A) $756.92 B) $437.50 C) $587.50 D) $563.66 @thesmartone or anyone please help !?

jimthompson5910 (jim_thompson5910):

I'm not 100% sure, but I think you solve it like this Annual Rate = 5% = 0.05 Monthly Rate = 0.05/12 = 0.00416666666667 Now multiply the monthly interest rate by the amount loaned out 0.00416666666667*105,000 = 437.5000000014 there is roundoff error and the amount is effectively $437.50